# DIPLI: Deep Image Prior Lucky Imaging for Blind Astronomical Image Restoration

### v1

@@ Abstract
Contemporary image restoration and super-resolution techniques effectively harness deep neural networks, markedly outperforming traditional methods. However, astrophotography presents unique challenges for deep learning due to limited training data. This work explores hybrid strategies, such as the Deep Image Prior (DIP) model, which facilitates blind training but is susceptible to overfitting, artifact generation, and instability when handling noisy images. We propose enhancements to the DIP model’s baseline performance through several advanced techniques. First, we refine the model to process multiple frames concurrently, employing the Back Projection method and the TVNet model. Next, we adopt a Markov approach incorporating Monte Carlo estimation, Langevin dynamics, and a variational input technique to achieve unbiased estimates with minimal variance and counteract overfitting effectively. Collectively, these modifications reduce the likelihood of noise learning and mitigate loss function fluctuations during training, enhancing result stability. We validated our algorithm across multiple image sets of astronomical and celestial objects, achieving performance that not only mitigates limitations of Lucky Imaging, a classical computer vision technique that remains a standard in astronomical image reconstruction but surpasses the original DIP model, state of the art transformer- and diffusion-based models, underscoring the significance of our improvements.
Index Terms:
High-resolution imaging, Image reconstruction, Astronomy, Telescopes, Digital photography.
@@
=-15pt
I Introduction
The imaging of astronomical objects is frequently associated with several challenges. In most cases, the resulting images are adversely affected by noise, blur, spatial distortion, and quality degradation, which arise from inherent limitations of the imaging system. Experienced astrophotographers can address some of these challenges through the utilization of advanced hardware and sophisticated post-processing methodologies. Nevertheless, super-resolution and image restoration—processes aimed at reconstructing high-quality (HQ) images from observed low-quality (LQ) counterparts—remain critical for amateur astrophotographers and small research groups.
Lucky Imaging. Deep learning (DL) has gained significant popularity in numerous applications of computer vision and computational photography  [1, 2], as well as astronomical imaging [3, 4]. However, classical analysis methods for astronomical imaging remain popular due to their reliability and interpretability features [5]. One of the most effective approaches is Lucky Imaging (LI) [6, 7, 8] that produces a single high-quality (HQ) image from a coherent sequence of lower-quality (LQ) frames. The method typically involves aligning each frame with a pivot with a subsequent averaging, as outlined in Algorithm 1. Through this process, LI can fast and efficiently eliminate noise and reduce atmospheric distortions on a frame-by-frame basis, which is why it is still actively employed in the data processing workflows of modern telescopes. However, LI also has notable limitations [9], including the need for a substantial number of source images, often on the order of several thousands. This requirement stems from the inverse relationship between the number of images averaged and the resulting noise level. In addition, its reliance on accurate motion estimation can hinder performance, as reported in [10]. Recent advances in motion estimation and robust frame selection [11] aim to overcome these limitations, yet challenges remain to handle highly dynamic scenes and faint objects.
Algorithm 1 Lucky Imaging Algorithm
0:  $\{x_{k}\}_{k=1}^{K}$: LQ frames; $w(\cdot,u_{k})$: spatial transform operator; quality metric $q(\cdot)$.
@@ I Introduction
1:  $x_{\mathrm{pivot}}\leftarrow\arg\max_{k}q(x_{k})$
2:  $u_{k}\leftarrow\mathcal{M}\bigl{(}x_{\mathrm{pivot}},x_{k}\bigr{)}\;\forall k$ motion compensation
3:  $x^{*}=\frac{1}{K}\sum_{k=1}^{K}w(x_{k},u_{k})$
4:  return  $x^{*}$
Our contribution. In this paper, we investigate the effectiveness of combining classical Lucky Imaging techniques with modern Unsupervised Neural Network Priors (UNNP), namely Deep Image Prior (DIP) to enhance observations of nearby space objects, such as planets, stars, and satellites, obtained with DSLR cameras mounted on optical telescopes. Despite the variety of deep learning-based methods, there is a noticeable gap in approaches that seamlessly integrate the strengths of both classical and machine learning paradigms [12]. Our proposed method, DIPLI, seeks to bridge this gap by combining the robustness of LI with the flexibility of DIP, creating a unified framework for blind image restoration that performs well in diverse scenarios. We demonstrate the applicability of our approach in this area and propose several improvements, namely prediction optical flows via unsupervised neural network TVNet [13], utilizing Stochastic Gradient Langevin Dynamics [14], and incorporating ideas from Back-Projections [15] method for preventing overfitting while training. Those improvements significantly increase reconstruction quality. On the one hand, the proposed method avoids the issues related to the number of frames required by Lucky Imaging, and on the other, it elegantly adapts machine learning techniques for astrophotography, an area that differs considerably from the classic domain of natural images on which most neural networks are trained.
We tested our method on both artificial and real data. We used advanced Deep Learning methods such as Diffusion Models and RVRT to assess the quality of the proposed algorithm.
II Related Work
Inverse Imaging Problems. Reconstructing a HQ scene from a set of LQ observations which is formulated in Eq.  1 is generally an ill-posed Inverse Imaging Problem (IIP)  [16, 17, 18] since the degradation model is generally a non-bijective function, meaning that there could be more than one feasible solution for the HQ:
@@ II Related Work
Deep Image Prior. The Untrained Neural Network Priors (UNNPs) framework  [23, 26, 27, 28, 25, 29, 30, 31] was originally proposed in [32], bridging the gap between traditional hand-crafted priors and DL approaches. This generalized framework (Fig.2) can accurately estimate clean samples from a single corrupted measurement without prior knowledge of the ground truth, outperforming conventional hand-crafted optimization. UNNPs leverage the rich image statistics captured by randomly initialized convolutional neural networks (CNNs), with the network weights serving as a parameterization of the restored image.
@@
where $y^{*}=G_{\theta^{*}}(z)$ is the reconstruction of $y$, and $z\sim\mathcal{N}(0,I)$ is a fixed random input noise (latent code).
@@
Bayesian inference. Consequently, selecting an appropriate coefficient for early stopping to mitigate this issue becomes necessary. Several heuristic techniques have been proposed to reduce this effect  [33, 34].
@@
To avoid the computation of this posterior average  [37], in standard Markov Chain Monte Carlo (MCMC)  [38] methods, the integral is replaced by a sample average of a Markov chain that converges to the true posterior average. However, convergence with MCMC techniques is significantly slower than Stochastic Gradient Descent (SGD) [39] for Deep Neural Networks (DNNs)[40]. Therefore, it is more convenient to use Stochastic Gradient Langevin Dynamics (SGLD) [41], which transforms SGD into an MCMC sampler by injecting noise into gradient updates.
III Method
In this section, we present DIPLI, an iterative approach to reconstructing a high-quality (HQ) scene $y$ from a set of $K$ distorted observations (low-quality or LQ frames) $X=(x_{1},x_{2},\dots,x_{K})$. Each LQ frame is modeled as:
$x_{k}=f_{k}(y)+\eta_{k},\quad\eta_{k}\sim\mathcal{N}(0,\sigma_{\eta}^{2}I),\quad\forall k\in\{1,\dots,K\},$ (6)
where $\eta_{k}$ generally contains multiple noise components. To simulate realistic background light pollution in artificial datasets, a mixture of Gaussian and Poisson noise is employed.
Degradation model. The mapping $f_{k}=d\circ h\circ\omega_{k}$ accounts for three major distortions:
@@ III Method
To choose $x_{\text{pivot}}$, the energy of the Laplacian [42] is computed for every frame, and the one maximizing this metric is selected:
@@
Iterative back-projection with a neural prior. Reconstruction of $y$ from $X$ proceeds via an iterative back-projection formulation [15]:
@@
Variational inference. To mitigate overfitting, a variational inference framework is adopted to target the minimum mean square error (MMSE) solution [37]. Instead of returning a single optimal $\theta^{*}$, the MMSE estimate is expressed as the expectation under the posterior distribution $p(\theta\mid X)$:
@@
Since direct sampling from $p(\theta\mid X)$ is intractable, stochastic gradient Langevin dynamics (SGLD) is employed:
@@
Figure 3: Comparison of DIPLI reconstruction quality with different number of received LQ frames.
@@
1:  $x_{\mathrm{pivot}}\leftarrow\arg\max_{x_{k}}q(x_{k})$
2:  $\omega_{k}\leftarrow\mathcal{F}\bigl{(}x_{\mathrm{pivot}},x_{k}\bigr{)}\;\forall k$
3:  $z\sim\mathcal{N}(0,I)$
4:  $\theta_{0}\sim p_{0}(\theta)$ random initial distribution
5:  for $n\in\{1,N\}$ do $z_{n}\sim\mathcal{N}(0,\sigma_{z}^{2}I)$
6:     $\xi_{n}\sim\mathcal{N}(0,\sigma_{\xi}^{2}I)$
@@
7:     $\theta_{n}\leftarrow\theta_{n-1}-\lambda_{n}\,\nabla_{\theta}\mathcal{L}\bigl{(}X,G_{\theta}(z+z_{n});\Omega\bigr{)}\Bigr{|}_{\theta=\theta_{n-1}}+\xi_{n}$
8:     if $n>n_{0}$ then
9:        $y^{*}\leftarrow y^{*}+G_{\theta_{n}}(z+z_{n})$
10:     end if
11:  end for
12:  return  $\frac{1}{N-n_{0}}y^{*}$
IV Experiments and Discussion
@@
To assess the quality of the proposed DIPLI model, we conducted two categories of experiments: one with artificial data and the other with real-world data. The real-world data consists of 15 videos capturing celestial objects, including the Moon, Mars, Saturn, Jupiter, and the Sun, provided at a resolution of $256\times 256$ pixels. These videos are sourced from private collection and open data. A notable challenge with real-world data is the absence of ground truth images for direct comparison, limiting validation methods to no reference metrics and human visual analysis. Our experiments employed the Laplace Energy metric for blind quality assessment, a standard measure used in both the DIPLI and Lucky Imaging algorithms during data registration and image ranking processes.
@@
The artificial dataset comprises artificially generated video sequences depicting the planetary images and data from Mars Exploration Rovers [44] both at $256\times 256$ pixels, alongside corresponding ground truth images at a resolution of $1024\times 1024$ pixels. The presence of the ground truth images enables to evaluate the model’s performance in terms of more trustworthy and interpretable reference-based metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [45], along with Neural Network-based metrics namely LPIPS [46], and DISTS [47]. Note that these ground truth images are not utilized in training the network, as our objective is to develop a blind super-resolution algorithm.
@@
For our evaluation, we used the DIPLI alongside modern Deep Learning models RVRT and Diffusion Model aiming to upscale input data by a factor of 4, denoise, and deblur. DIPLI was trained for 6500 iterations, averaging the reconstructions from the last 500 iterations to obtain the final result as a Monte Carlo estimation. Moreover, we defined the crucial parameter: the optimal number of source low-quality (LQ) images provided to DIPLI, RVRT, and DiffIR2VR-Zero is $i=11$, where a pivot image is preceded by 5 coherently previous images and followed by 5 coherently upcoming images. To obtain this optimum, we tested several different values ranging from 1 to 100 and displayed the results in Figure 3. As can be seen, the most accurate reconstruction is achieved for 10 frames.
IV-A Motion Compensation
Generally, there are numerous ways to perform motion compensation. The simplest way is to find a vector shift between two images. This method is fast and remains effective for the majority of applications where solid spatial motions occur caused by the camera shaking. In most cases, a single vector shift is enough since no other more complex motion occurs. Unfortunately, when the imaging subject is in motion, or heat-induced distortions are present, a simple shift is no longer sufficient. In such cases, optical flow becomes necessary. Optical flow extends the concept of a single vector shift to a vector field, assigning a motion vector to each pixel in the image to account for its individual movement.
@@ IV-A Motion Compensation
There are several possible ways to construct an optical flow with probably the most common way, the TV-L1 [48]. However, noisy images pose a significant challenge in constructing optical flows since they corrupt source images and create an obstacle to a true map of the pixel’s transpositions. To address such challenges, neural network-based approaches, like TVNet, have been developed and implemented. Without training, TVNet mirrors the behavior of TV-L1, but with additional unsupervised tuning, it can overcome noise. Below, in the Figure 4 we provide results of corresponding experimental evidence aiming to evaluate different motion compensation methods: GRAVITY, directly calculating a spatial shift between two centers of masses in the images, classical Iterative Lucas-Kanade (ILK) method [49], basic TV-L1 method, and two neural network methods RAFT-L [50] and TVNet that we used in the proposed model. In order to assess the method performance we apply them for a pair of images to find a transformation that matches one to the other. Next, we apply the obtained transformation to the corresponding image and compare the result with the other image in terms of a common Mean Average Error (MAE) metric to determine the similarity.
@@
Figure 5: Comparison of DIPLI reconstruction quality with different values of the SGLD strength coefficient $\sigma_{\xi}$.
Figure 6: Synthetic data reconstructions. Insets are added to guide the eye.
IV-B SGLD strength
As mentioned above, UNNP models tend to overfit and memorize the noise. To avoid this, Bayesian Inference offers a special MCMC strategy called SGLD. It implies forceful regularization with additional noise applied to the learnable parameters on every optimization step to prevent overfitting as a process of stagnating in one solution and intermediate averaging to withstand the noise influence. Here, we provide extensive experiments to confirm the advantage of the SGLD strategy for overfit reduction. However, SGLD indeed delivers excellent results, and it is crucial to strike a balance in the regularization value that is the strength of the additional noise. If it is too low, the impact of the strategy becomes barely noticeable, while excessively large value not only prevent overfitting but also hinder the learning process altogether. Experiments shown in the Figure 5 indicate that the best noise variation coefficient value is $\sigma_{\xi}=0.0025$ (same as the chosen learning rate for the experiments).
In addition to SGLD strategy, we use an additional noise while generating a latent vector $z$ from the initial distribution. However, it was stated in the original paper of Bayesian DIP [37] that this additional noise does not bring any advantage. We keep it for the sake of smoothness in a latent space and potential resistance to adversarial perturbations in latent space  [51], according to the original DIP paper [52].
IV-C Synthetic Data
Finally, this section reports the main experimental results starting with synthetic data. The synthetic dataset is created with a degradation model described above. Namely, we take one ground truth picture at high resolution and gradually corrupt it with noise and a variety of spatial deformations simulating realistic process of video shooting of a celestial object.
The results depicted in Figure 6 and Tables I,  II show that the proposed DIPLI approach demonstrates the best performance in terms of the most metrics and achieves a fair balance between reconstruction details and noise reduction compared to the aforementioned alternative methods. Regarding the other baselines, a basic configuration of RVRT shows a tendency to overfit, which can be seen as a finely dispersed pattern in the reconstruction. The RVRT+ model with a preliminary denoising process (utilizing the same model with different weights) avoids straightforward noise reconstruction. However, it does not show a sufficient level of image improvement compared to the original LQ observation. As for DiffIR2VR-Zero, known for high-quality image generation, it indeed performs comparably well in our task, producing sharp and visually appealing images. However, it often smooths out certain areas, removing critical details present in the original data and generally tend to hallucinate textures and details. This propensity to prioritize aesthetics over accuracy is problematic in fields such as astrophotography, where preserving fine details is essential to prevent misinformation and unreliable results.
@@ IV-C Synthetic Data
TABLE I: PSNR and SSIM metrics of the observation and ground truth compared to the corresponding values for synthetic data reconstructions obtained by DIPLI, RVRT, and DiffIR2VR-Zero, respectively.

### v2

@@ Abstract
Modern image restoration and super-resolution methods utilize deep learning due to its superior performance compared to traditional algorithms. However, deep learning typically requires large training datasets, which are rarely available in astrophotography. Deep Image Prior (DIP) bypasses this constraint by performing blind training on a single image. Although effective in some cases, DIP often suffers from overfitting, artifact generation, and instability. To overcome these issues and improve general performance, this work proposes DIPLI - a framework that shifts from single-frame to multi-frame training using the Back Projection technique, combined with optical flow estimation via the TVNet model, and replaces deterministic predictions with unbiased Monte Carlo estimation obtained through Langevin dynamics. A comprehensive evaluation compares the method against Lucky Imaging, a classical computer vision technique still widely used in astronomical image reconstruction, DIP, the transformer-based model RVRT, and the diffusion-based model DiffIR2VR-Zero. Experiments on synthetic datasets demonstrate consistent improvements, with the method outperforming baselines for SSIM, PSNR, LPIPS, and DISTS metrics in the majority of cases. In addition to superior reconstruction quality, the model also requires far fewer input images than Lucky Imaging and is less prone to overfitting or artifact generation. Evaluation on real-world astronomical data, where domain shifts typically hinder generalization, shows that the method maintains high reconstruction quality, confirming practical robustness.
1 Introduction
Imaging astronomical objects presents persistent challenges due to atmospheric turbulence, sensor noise, optical blur, and spatial distortions. These degradations stem from fundamental limitations in imaging hardware and observation conditions. While advanced equipment can mitigate some of these effects, software-based techniques, particularly super-resolution and image restoration, remain essential for reconstructing high-quality (HQ) images from low-quality (LQ) observations. Such methods are crucial not only for amateur astrophotographers and small research groups with limited resources but also for professional teams seeking to enhance data acquired under suboptimal conditions.
Among classical methods valued for their reliability and interpretability [1], Lucky Imaging (LI) [2, 3, 4] continues to be one of the most commonly used. LI reconstructs an HQ image from an unordered set of LQ frames by selecting a pivot frame, aligning all other frames to it via motion compensation, and subsequently averaging, as outlined in Algorithm 1. This method effectively suppresses noise and atmospheric distortions and remains actively used in telescope data pipelines due to its simplicity and reliability. However, the method also has notable limitations [5], including the need for a substantial number of source images, typically in the range of several thousand, to effectively suppress noise. In addition, its reliance on accurate pivot selection and motion estimation can hinder performance, as reported in [6]. While recent advances in motion compensation and frame selection [7] have improved its effectiveness, LI still struggles with highly dynamic scenes and faint celestial objects.
Algorithm 1 Lucky Imaging Algorithm
@@ 1 Introduction
[…]
4:return $x^{*}$
Deep Learning (DL) has demonstrated remarkable success in a variety of image restoration scenarios [8, 9, 10, 11]. However, most DL models require extensive training data, which is typically unavailable in astrophotography. While pretrained models offer a potential solution, the significant domain gap between astronomical and natural images, on which most models are trained, makes them ineffective. Consequently, models capable of learning from limited data become necessary. One such approach is Deep Image Prior (DIP) [12], which performs unsupervised optimization on a single image. Despite its appeal, DIP is highly sensitive, prone to overfitting, and lacks robustness in real-world conditions.
To address the limitations of existing methods, this work introduces a unified framework DIPLI that combines the robustness of Lucky Imaging with the flexibility of Deep Image Prior. The proposed model operates effectively with only 7–11 input frames by integrating Back Projection with optical flow estimated by the unsupervised TVNet model [13]. To reduce overfitting and improve robustness, DIPLI replaces deterministic inference with Monte Carlo averaging using simplified Stochastic Gradient Langevin Dynamics (SGLD) [14] for sampling. This hybrid approach bridges classical and learning-based paradigms [15], achieving high-quality image reconstruction while minimizing data requirements and sensitivity to noise.
This work offers the following key contributions:
  1. 1.
    The model extends DIP to operate on multiple frames using the Back Projection technique and optical flow estimated by TVNet.
  2. 2.
    The approach employs SGLD for sampling in Monte Carlo estimation.
  3. 3.
    Comprehensive evaluation includes extensive benchmarking against classical Lucky Imaging, original DIP, and state-of-the-art deep learning models, such as RVRT [16] and DiffIR2VR-Zero [17], on both synthetic and real astronomical datasets, using PSNR, SSIM [18], LPIPS [19], and DISTS [20] as performance metrics.
  4. 4.
    Exhaustive ablation studies the quality of motion compensation, influence of the number of input frames, and SGLD noise strength to assess their impact on reconstruction quality.
This paper is organized as follows: Section 2 reviews related work in astronomical imaging, classical restoration methods, and neural priors. Section 3 details the proposed DIPLI method. Section 4 presents experimental results, comparative evaluations, and ablation studies along with exhaustive analysis. Finally, section 5 concludes with a summary and potential directions for future research.
2 Related Work
Inverse Imaging Problems.
Reconstructing a HQ scene from a set of LQ observations which is formulated in Eq. 1 is generally an ill-posed Inverse Imaging Problem (IIP) [21, 22, 23] since the degradation model (forward operator) is generally a non-bijective function, meaning that there could be more than one feasible solution for the HQ:
$\text{LQ}=\text{Degradation Model}(\text{HQ}).$ (1)
While the form of the degradation model depends on the specific problem, a unified degradation framework is described in detail in Section 3. Some methods, such as Lucky Imaging, avoid directly using the inverse of a degradation model (backward operator). This approach ensures determinism and uniqueness of the result obtained. However, it does not consider the complete solution domain, and therefore does not guarantee global optimality. On the other hand, methods that seek to examine multiple candidates in the space of potential solutions P face enormous computational challenges. In the case of the image domain, this space, or more accurately, distribution $P(\text{HQ}|\text{LQ})$ exhibits unpredictably complex statistics properties and is generally intractable.
@@ 2 Related Work
Deep Image Prior.
The Untrained Neural Network Priors (UNNPs) framework [28, 31, 32, 33, 30, 34, 35, 36] was originally proposed in [12], bridging the gap between traditional hand-crafted priors and DL approaches. This generalized framework (Fig.2) can accurately estimate clean samples from a single corrupted measurement without prior knowledge of the ground truth, outperforming conventional hand-crafted optimization. UNNPs leverage the rich image statistics captured by randomly initialized convolutional neural networks (CNNs), with the network weights serving as a parameterization of the restored image.
@@
where $y^{*}=G_{\theta^{*}}(z)$ is the reconstruction of $y$, $f$ is a degradation function or an imaging system model, and $z\sim\mathcal{N}(0,I)$ is a fixed random input noise (latent code).
@@
Bayesian inference.
Consequently, selecting an appropriate coefficient for early stopping to mitigate this issue becomes necessary. Several heuristic techniques have been proposed to reduce this effect [37, 38].
@@
To avoid the computation of this posterior average [41], in standard Markov Chain Monte Carlo (MCMC) [42] methods, the integral is replaced by a sample average of a Markov chain that converges to the true posterior average. However, convergence with MCMC techniques is significantly slower than Stochastic Gradient Descent (SGD) [43] for Deep Neural Networks (DNNs)[44]. Therefore, it is more convenient to use SSGLD [45], which transforms SGD into an MCMC sampler by injecting noise into gradient updates.
3 Method
In this section, we present DIPLI, an iterative approach to reconstructing a high-quality (HQ) scene $y^{*}$ from a set of $K$ distorted observations (low-quality or LQ frames) $X=(x_{1},x_{2},\dots,x_{K})$. Each LQ frame is modeled as:
$x_{k}=f_{k}(y)+\eta_{k},\quad\forall k\in\{1,\dots,K\},$ (6)
where $\eta_{k}$ can be modeled as Gaussian noise $\eta_{k}\sim\mathcal{N}(0,\sigma_{\eta}^{2}I)$ or consist of multiple components. To simulate realistic background light pollution in artificial datasets, a mixture of Gaussian and Poisson noise is employed.
Degradation model.
The mapping $f_{k}=d\circ h\circ\omega_{k}$ accounts for three major distortions:
@@ 3 Method
To choose $x_{\text{pivot}}$, the energy of the Laplacian [46] $q(x_{k})$ is computed for every frame $x_{k}$, and the one maximizing this metric is selected:
@@
Iterative back-projection with a neural prior.
Reconstruction of $y$ from $X$ proceeds via an iterative back-projection formulation [48]:
@@
Variational inference.
To mitigate overfitting, a variational inference framework is adopted to target the minimum mean square error (MMSE) solution [41]. Instead of returning a single optimal $\theta^{*}$, the MMSE estimate is expressed as the expectation under the posterior distribution $p(\theta\mid X)$:
@@
Since direct sampling from $p(\theta\mid X)$ is intractable, SGLD is employed:
@@
[…]
5:for $n\in\{1,N\}$ do
6:  $z_{n}\sim\mathcal{N}(0,\sigma_{z}^{2}I)$
7:  $\xi_{n}\sim\mathcal{N}(0,\sigma_{\xi}^{2}I)$
@@
8:  $\theta_{n}\leftarrow\theta_{n-1}-\lambda_{n}\,\nabla_{\theta}\mathcal{L}\bigl(X,G_{\theta}(z+z_{n});\Omega\bigr)\Bigr|_{\theta=\theta_{n-1}}+\xi_{n}$
9:  if $n>n_{0}$ then
10:   $y^{*}\leftarrow y^{*}+G_{\theta_{n}}(z+z_{n})$
11:  end if
12:end for
13:return $\frac{1}{N-n_{0}}y^{*}$
4 Experiments and Discussion
4.1 Data availability
All data generated or analysed during this study are included in this published article. The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
4.2 Dataset description
The quality of the proposed DIPLI model was evaluated using two categories of experiments: one based on artificially generated data and the other on real-world data.
@@ 4.2 Dataset description
The real-world data consists of 15 videos capturing celestial objects, including the Moon, Mars, Saturn, Jupiter, and the Sun, provided at a resolution of $256\times 256$ pixels. These videos are sourced from private collection and publicly available data. A primary challenge with real-world data is the absence of ground truth images for direct comparison, which restricts validation methods to no-reference quality metrics and human visual inspection. The energy of Laplacian [49] was used for blind quality assessment, as it is a standard metric used in Lucky Imaging algorithms during data registration and image ranking processes.
@@
The artificial dataset comprises artificially generated video sequences depicting the planetary images and data from Mars Exploration Rovers [50] both at $256\times 256$ pixels, alongside corresponding ground truth images at a resolution of $1024\times 1024$ pixels. The presence of the ground truth images enables to evaluate the model’s performance in terms of more trustworthy and interpretable reference-based metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [18], along with Neural Network-based metrics namely LPIPS [19], and DISTS [20]. It should be noted that the ground-truth images are not used for network training, as the objective is to develop a blind super-resolution algorithm.
The artificial dataset was generated using the degradation model described in Section 3. Specifically, a high-resolution ground-truth image was progressively corrupted with noise and various spatial distortions to simulate the realistic process of video acquisition of a celestial object.
Figure 3: Comparison of DIPLI reconstruction quality with different number of received LQ frames. While multiple frames provide complementary information, excessive inputs cause the model to fit inter-frame variations, reducing precision. Optimal performance occurs with approximately 7–13 frames.
4.3 Experimental Setup
The backbone of DIPLI is a four-stage U-Net with $128$ feature channels per stage, instance normalization, and symmetric skip connections between encoder and decoder paths. Each stage consists of two $3\times 3$ convolutions with ReLU activations and dropout, with downsampling performed via pooling and upsampling via bilinear interpolation followed by feature projection. The output is produced through a multi-step upsampling head that restores spatial resolution by a selected scaling factor $s\in\{2,4,8\}$ and applies a final convolution with sigmoid activation.
@@ 4.3 Experimental Setup
The evaluation compares DIPLI with modern deep learning models RVRT and DiffIR2VR-Zero, targeting super-resolution by the factor $4$, denoising, and deblurring. DIPLI was trained for 6500 iterations, averaging the reconstructions from the last 500 iterations to obtain the final result as a Monte Carlo estimation as described in Algorithm 2. The optimal number of source low-quality (LQ) images provided to DIPLI, RVRT, and DiffIR2VR-Zero was determined to be $K=11$, corresponding to a pivot image preceded by five temporally adjacent frames and followed by five subsequent frames. This value was obtained by testing a range of input sizes from 1 to 100, with results shown in Figure 3, where the highest reconstruction accuracy is observed for ten neighboring frames.
4.4 Experiments and Results
Motion Compensation.
Numerous approaches exist for motion compensation. The simplest estimates a single shift vector between two images, offering high speed and sufficient accuracy for scenarios dominated by uniform spatial motion, such as camera shake. In most cases, a single vector this single vector adequately models the displacement. However, when the subject moves or heat-induced distortions occur, a global shift becomes insufficient. Optical flow addresses this limitation by extending the single-vector concept to a dense vector field, assigning an individual motion vector to each pixel to capture localized movements.
@@ 4.4 Experiments and Results
There are several possible ways to construct an optical flow with probably the most common way, the TV-L1 [51]. Unfortunately, noisy images significantly hinder optical flow estimation by corrupting the source data and obscuring the true pixel displacement map. To address such challenges, neural network-based approaches, such as TVNet, have been developed and implemented. Without training, TVNet mirrors the behavior of TV-L1, but with additional unsupervised tuning, it can overcome noise. Figure 4 presents experimental results comparing different motion compensation methods: GRAVITY, directly calculating a spatial shift between two centers of masses in the images, classical Iterative Lucas-Kanade (ILK) method [52], basic TV-L1 method, and two neural network methods RAFT-L [53] and TVNet. To assess performance, each method estimates a transformation between an image pair (source, target). The transform warps the source to the target, and similarity is quantified by the mean absolute error (MAE) between the warped source and the target; lower MAE indicates better alignment.
@@
SGLD strength.
UNNP models are prone to overfitting, often memorizing image noise without explicit regularization. To mitigate this, the method adopts Bayesian inference via SGLD MCMC [41]. SGLD injects Gaussian noise into the parameter updates at each step, promotes posterior exploration, and prevents collapse to a single overfitted solution; averaging late-iteration samples further improves robustness to observation noise.
Regularization strength (i.e., the SGLD noise scale) requires careful tuning: too small has negligible effect, whereas too large impedes learning. Experiments in Figure 5 indicate an optimal noise standard deviation of $\sigma_{\xi}=0.0025$ (equal to the learning rate).
In addition to SGLD, a small perturbation is applied when sampling the latent vector $z$ from the prior. Although Bayesian DIP reports limited benefit from this perturbation [41], the pipeline retains it to promote latent-space smoothness and potential robustness to adversarial latent perturbations, consistent with DIP and randomized smoothing [54, 55].
Figure 5: Comparison of DIPLI reconstruction quality with different values of the SGLD strength coefficient $\sigma_{\xi}$. Following the approach of [41], the results indicate a tradeoff between noise impedance and the number of iterations required for convergence. Higher $\sigma_{\xi}$ values increase the convergence time. For astronomical images in this dataset, the optimal setting was $\sigma_{\xi}=0.0025$ with 6500 iterations.
Synthetic Data.
Figure 6 and Table 1 indicate that DIPLI achieves the best overall performance across most metrics and provides a favorable trade-off between detail preservation and noise suppression relative to the compared methods.
The baseline RVRT configuration exhibits overfitting, visible as fine-grained speckle in the reconstructions. RVRT+ applies preliminary denoising using the same architecture with different weights to reduce direct noise reproduction but yields only modest improvement over the original low-quality (LQ) input. DiffIR2VR-Zero, while producing sharp and visually appealing images, frequently smooths regions and removes salient details from the data. Diffusion-based model also tends to hallucinate textures. Such emphasis on aesthetics over fidelity is problematic in astrophotography, where preserving fine structure is essential to avoid misleading interpretations.
A direct, controlled comparison to DIP is nontrivial because DIP relies on early stopping to prevent overfitting, and the optimal stopping iteration is data-dependent and unknown without ground truth. For the synthetic experiments, a single stopping iteration was selected by scanning iteration counts over the dataset and choosing the value that maximized average PSNR and SSIM; this fixed iteration was then applied uniformly to all DIP runs, consistent with DIPLI’s fixed-iteration regime. On real data, the absence of references makes such tuning infeasible, leaving the DIP stopping point indeterminate and precluding a fair like-for-like comparison.
Figure 6: Synthetic-data reconstructions. The LQ image pivot is shown on the left, and the remaining images present the corresponding HQ restorations. Insets are included to highlight specific details.